# Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability

### v1

@@ 2.4 Theoretical Results
where the first inequality holds due to Lemma 2.17 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.16. The term $G(0)$ can be further bounded by
$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}$
$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (2.7)
3 KL-Regularized Contextual Dueling Bandits
In this section, we extend our algorithm to KL-regularized dueling bandit setting, where the learner receives preference feedback instead of absolute signals. We formally introduce setting and extended algorithm. Except the objective, the core setting follows Zhu et al. (2023); Zhan et al. (2023). Our objective follows Xiong et al. (2024); Zhao et al. (2024).
3.1 Problem Setup
We still [^4] We overload some notations in Section 2 by their dueling counterparts for notational simplicity. consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$. The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ and 00 for the opposite. Given a state $s$ and two corresponding action $a^{1}$ and $a^{2}$, let $y$ be the preference label such that $y=1$ indicates $a^{1}\succ a^{2}$ and $y=0$ for $a^{2}\succ a^{1}$. In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function. The objective here identical to (2.1) and thus the optimal policy $\pi^{*}$ also follows the closed form in (2.2). Our goal is still to find an $\epsilon$-optimal policy. To control the complexity of the function class $\mathcal{F}$, we assume that Assumption 2.1 still holds here.
Concentrability
Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy, and therefore requires similar $D^{2}$-divergence and corresponding policy coverage conditions. However, in dueling bandit, we are unable to observe and estimate the absolute value of rewards. Therefore, intuitively, the best estimation $f$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $f(s,a^{1})-f(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$. This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $f(s,a)-b(s)\approx r(s,a)$ on the offline data. Consequently, it is natural to define the the following modified $D^{2}$-divergence for characterizing how the error $f(s,a)-b(s)-r(s,a)$ generalize to unseen samples.
Definition 3.1.
Given a class of functions $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
$\sup_{f,g\in\mathcal{F}}\inf_{b\in\mathcal{B}}\frac{\big{(}f(s,a)-g(s,a)-b(s)\big{)}^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime})]}.$
A similar definition has been introduced in Zhao et al. (2024, Definition 2.6), which underpins the following two assumptions characterizing the coverage ability of $\pi^{\mathsf{ref}}$.
Equipped with the overloaded $D^{2}$-divergence definition, we are now able to define all-policy concentrability and single concentrability similar as in Section 2.
Assumption 3.2 (All-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
Assumption 3.3 (Single-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
Remark 3.4.
Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits. Specifically, Huang et al. (2024) defined an $L_{\infty}$-concentrability coefficient $\mathcal{C}_{\infty}^{\pi^{*}}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\big{(}\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\big{)}$. Song et al. (2024) introduced a stronger assumption that $\mathcal{C}_{\infty}^{\pi}$ is bounded for all the policy $\pi$ in the KL-ball surrounding $\pi^{*}$. In our Definition 3.1, instead of simply considering the ratio of probability measure, we introduce a more refined eluder-type uncertainty term which also relies on the structure of the model class. As a result, $D_{\pi^{*}}$ can be small even if $L_{\infty}$-concentrability coefficient of $\pi^{*}$ is extremely large as long as the reward model generalizes well after training on the offline dataset. There has also been similar notions of partial coverage for the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024), which fused the information of the transition kernel class to the density-ratio-based formulation.
3.2 Algorithm
@@
We elucidate KL-PCDB for offline KL-regularized contextual dueling bandits, whose pseudocode is summarized in Algorithm 2. KL-PCDB first estimate the ground truth function $f^{*}$ on offline dataset with maximum likelihood estimator (MLE) to estimate a function ${\bar{f}}\in\mathcal{F}$. After that, analogous to Algorithm 1, we adopt the principle of pessimism in the face of uncertainty. Specifically, we define the penalty term
$\Gamma_{n}(s,a)=\beta\sqrt{D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (3.1)
@@
$\beta^{2}=128\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (3.2)
and then subtract it from the MLE ${\bar{f}}$ to obtain a pessimistic estimator ${\widehat{f}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
@@
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
1:  Compute the maximum likelihood estimator of the reward function
${\bar{f}}=\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}f(s_{i},a^{1}_{i})-f(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}f(s_{i},a^{2}_{i})-f(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}$
2:  Let ${\widehat{f}}(s,a)={\bar{f}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in 3.1
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
3.3 Main Results
In this subsection, we provide our main results for KL-regularized contextual dueling bandits. The sample complexity upper bound of KL-PCDB is presented in the following theorem.
Theorem 3.5.
Under Assumption 3.3, if we set $\Gamma_{n}$ according to Lemma 3.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 2 to be $\epsilon$-optimal.
Remark 3.6.
@@ 3.3 Main Results
Theorem 3.7.
For any $\epsilon\in(0,1)$, $\eta>0$, and offline RL algorithm Alg, there is a KL-regularized contextual dueling bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.

### v2

@@ Appendix C Missing Proof from Section 3
Consider our estimation ${\bar{g}}$ which approximates the ground truth reward function ${g^{*}}$, we know that
$\widehat{\pi}=\mathop{\mathrm{argmax}}_{\pi\in{\mathcal{S}}\to\Delta(\mathcal{A})}\Big{\{}\mathbb{E}_{(s,a)\sim\rho\times\pi}[{\bar{g}}(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]}\Big{\}}.$
We have the following sub-optimality decomposition
$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
$=\mathbb{E}_{s\sim\rho}\big{[}H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}})-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}\right\rangle\big{]}$
$=\mathbb{E}_{s\sim\rho}\big{[}H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}})-\left\langle\nabla H^{*}_{s}({\bar{g}}),{g^{*}}-{\bar{g}}\right\rangle\big{]}$
$=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}})],$
where $\widetilde{g}=\gamma{g^{*}}+(1-\gamma){\bar{g}}$ and $\gamma\in[0,1]$ and the last equation holds due to Taylor’s expansion. Now, for any $\delta\in(0,1)$ and $\epsilon_{c}>0$, with probability at least $1-\delta$
$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}})]$
$\leq\alpha^{-1}\eta\mathbb{E}_{s\sim\rho}\Big{[}({g^{*}}-{\bar{g}})^{\top}\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)\big{)}({g^{*}}-{\bar{g}})\Big{]}$
$=\alpha^{-1}\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{\mathsf{ref}}}\big{[}\big{(}{g^{*}}(s,a)-{\bar{g}}\big{(}s,a))^{2}\big{]}$
$\leq\alpha^{-1}\eta\bigg{(}\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c}\bigg{)},$
where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (B.1). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
C.2 Proof of Theorem 3.4
We first provide the following lemma that gives the close form of optimal policy under $\chi^{2}$-divergence regularization.
Lemma C.5 (Huang et al. (2025, Lemma G.2)).
Let $\pi^{*}$ be the optimal policy of $\chi^{2}$-divergence regularized objective with reward function $r$, then $\pi^{*}$ has the closed form
$\pi^{*}(\cdot)=\pi^{\mathsf{ref}}(\cdot)\max\big{\{}0,\eta(r(\cdot)-\lambda)\big{\}},\text{ where }\sum_{a\in\mathcal{A}}\pi^{*}_{f\mathrm{div}}(a)=1.$
By Proposition C.2, $\pi^{*}_{f\mathrm{div}}=\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A})}J_{f\mathrm{div}}(\pi)$ is unique. The sub-optimality gap for $f$-divergence is consequently defined as
$\mathrm{SubOpt}_{f\mathrm{div}}(\cdot)\coloneqq\mathrm{SubOpt}_{f\mathrm{div}}(\cdot;\mathcal{A},r,\pi^{\mathsf{ref}})=J_{f\mathrm{div}}(\pi^{*}_{f\mathrm{div}})-J_{f\mathrm{div}}(\cdot).$ (C.2)
Now we are ready to prove Theorem 3.4.
Proof of Theorem 3.4.
We still consider the family of contextual bandits $\mathrm{CB}_{\mathcal{G}}$ given by (B.3). We, still, aim to prove the following statement. Fixing any $S\geq 32\log 2$, $\eta>4\log 2$ and $\alpha$, we set $f(x)\coloneqq\alpha(x-1)^{2}/2$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n$ sufficiently large, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
$\mathrm{SubOpt}_{f}(\widehat{\pi};\mathrm{inst})\gtrsim\alpha^{-1}\eta Sn^{-1}.$ (C.3)
Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (C.3) yields the desired bound.
[…]
$\mathcal{G}=\{r_{v}(s,-1)=1/2+v_{s}\delta,r_{v^{\prime}}(s,+1)=1/2+v_{s}\delta,\ \forall s\in{\mathcal{S}}|v\in\mathcal{V}\},$
where $\delta\in(0,\eta^{-1}\alpha]$ is to be specified later. Fix some context $s$ and $v_{1}\neq v_{2}$ different at entry $s$ and corresponding reward $r_{1}$ and $r_{2}$. Without loss of generality, we assume $r_{1}(s,\cdot)=(1/2+\delta,1/2-\delta)$ and $r_{2}(s,\cdot)=(1/2-\delta,1/2+\delta)$. Then direct calculation implies that
$\pi_{1}^{*}(\cdot|s)=\frac{1}{2}\max\{0,\eta\alpha^{-1}(r_{1}(s,\cdot)-{\lambda})\}=0.5\eta\alpha^{-1}(r_{1}(s,\cdot)-{\lambda}),$
$\pi_{2}^{*}(\cdot|s)=\frac{1}{2}\max\{0,\eta\alpha^{-1}(r_{2}(s,\cdot)-{\lambda})\}=0.5\eta\alpha^{-1}(r_{2}(s,\cdot)-{\lambda}),$
where $\lambda=0.5-\eta^{-1}\alpha$. Note that $2{\chi^{2}}(\mu\|\nu)+1=\sum_{a\in\mathcal{A}}[\mu(a)]^{2}/\nu(a)$ and ${\chi^{2}}=D_{f}$, we obtain that $\forall\ \widehat{\pi}$,
$\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{1})+\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{2})$ (C.4)
$\quad=\langle{r_{1}(s,\cdot)},{\pi^{*}_{1}(\cdot|s)}\rangle+\langle{r_{2}(s,\cdot)},{\pi^{*}_{2}(\cdot|s)}\rangle-\overbrace{\langle{r_{1}(s,\cdot)+r_{2}(s,\cdot)},{\widehat{\pi}(\cdot|s)}\rangle}^{=1}+\overbrace{2\eta^{-1}\alpha{\chi^{2}}(\widehat{\pi}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))}^{\geq 0}$
$\quad\quad-\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{1}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))-\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{2}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$
$\quad\geq 2\langle{r_{1}(s,\cdot)},{\pi^{*}_{1}(\cdot|s)}\rangle-1-2\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{1}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$
$\quad=1+\frac{2\eta\delta^{2}}{\alpha}-1-\frac{\eta\delta^{2}}{\alpha}=\frac{\eta\delta^{2}}{\alpha}.$ (C.5)
Now we take expectation over all possible contexts and recall that $\|v-v^{\prime}\|_{1}\geq S/2$ for $v\neq v^{\prime}$, we know that for any $r_{1}\neq r_{2}\in\mathcal{G}$
$\mathrm{SubOpt}(\widehat{\pi};r_{1})+\mathrm{SubOpt}(\widehat{\pi};r_{2})\geq\frac{\eta\delta^{2}}{2\alpha}$
Given any mean reward function $r\in\mathcal{G}$, let $P_{r}$ be the distribution of $(s,a,\mathtt{r})$ when $s\sim\rho$, $a\sim\pi^{\mathsf{ref}}(\cdot|s)$, and $\mathtt{r}\sim\mathsf{Bern}(r(s,a))$. Suppose $P_{\mathcal{D}_{r}}$ is the distribution of the dataset given mean reward function $r$, then ${\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)=n{\mathsf{KL}}\left(P_{{r_{1}}}\|P_{{r_{2}}}\right)$ for any pair of $r_{1},r_{2}\in\mathcal{G}$. Now we invoke Fano’s inequality (Lemma F.2) to obtain
$\inf_{\pi}\sup_{\mathrm{inst}\in\mathrm{CB}_{\mathcal{G}}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})$ $\geq\frac{\eta\delta^{2}}{4\alpha}\Bigg{(}1-\frac{\max_{r_{1}\neq r_{2}\in\mathcal{G}}{\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)+\log 2}{\log|\mathcal{G}|}\Bigg{)}$
$\geq\frac{\eta\delta^{2}}{4\alpha}\Bigg{(}1-\frac{64n\delta^{2}+8\log 2}{S}\Bigg{)},$
where the second inequality holds due to $\mathsf{KL}(\mathsf{Bern}(p)\|\mathsf{Bern}(q))\leq(p-q)^{2}/[q(1-q)]$. Let $\delta=16\sqrt{\alpha\eta^{-1}n^{-1}}$, then we obtain that for all $\pi$ we have
$\sup_{\mathrm{inst}\in\mathrm{CB}_{\mathcal{G}}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$
which finishes the proof in that $\log_{2}|\mathcal{G}|=S$. ∎
Appendix D Missing Details from Section 4
In this section, we provide additional details that has been omitted in Section 4.
D.1 Algorithms for KL-regularized Contextual Dueling Bandits
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class ${\mathcal{G}}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
1:  Compute the maximum likelihood estimator of the reward function
${\bar{g}}=\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}g(s_{i},a^{1}_{i})-g(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}g(s_{i},a^{2}_{i})-g(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}$
2:  Let ${\widehat{g}}(s,a)={\bar{g}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in D.1
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}$
@@ D.1 Algorithms for KL-regularized Contextual Dueling Bandits
We elucidate KL-PCDB for offline KL-regularized contextual dueling bandits, whose pseudocode is summarized in Algorithm 3. KL-PCDB first estimate the ground truth function ${g^{*}}$ on offline dataset with maximum likelihood estimator (MLE) to estimate a function ${\bar{g}}\in{\mathcal{G}}$. After that, analogous to Algorithm 1, we adopt the principle of pessimism in the face of uncertainty. Specifically, we define the penalty term
$\Gamma_{n}(s,a)=\beta\sqrt{D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (D.1)
@@
$\beta^{2}=128\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (D.2)
and then subtract it from the MLE ${\bar{g}}$ to obtain a pessimistic estimator ${\widehat{g}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{g}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
@@
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}.$
We provide the following theoretical guarantees for Algorithm 3.
Theorem D.1.
Under Assumption 4.3, if we set $\Gamma_{n}$ according to D.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\epsilon^{-1}\big{)}$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 3 to be $\epsilon$-optimal.
Remark D.2.
@@
Theorem D.3.
For any sufficiently small $\epsilon\in(0,1),\eta>0$, $21\leq C^{*}\leq\exp(\eta/2)/$ 2, and any algorithm Alg, there is a KL-regularized contextual dueling bandit instance with single-policy concentrability $D^{2}_{\pi^{*}}\leq 2C^{*}$ such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}/{\epsilon},{\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}(C^{*})^{2}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Remark D.4.
Theorem D.3 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandits requires at least $\Omega(\eta D^{2}_{\pi^{*}}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy, which matches the sample complexity upper bound in Theorem D.1, indicating that KL-PCB is nearly optimal.
D.2 Algorithm and Results for $f$-divergence Regularized Contextual Dueling Bandit
Algorithm 4 Offline $f$-Divergence Regularized Contextual Dueling Bandits ($f$-CDB)
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class ${\mathcal{G}}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
1:  Compute the maximum likelihood estimator of the reward function
${\bar{g}}=\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}g(s_{i},a^{1}_{i})-g(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}g(s_{i},a^{2}_{i})-g(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}.$
2:  Compute the optimal policy with respect to reward ${\bar{g}}$
$\widehat{\pi}(\cdot|s)\leftarrow\mathop{\mathrm{argmax}}_{\pi(\cdot|s)\in\Delta(\mathcal{A})}\sum_{a\in\mathcal{A}}\pi(a|s){\bar{g}}(s,a)+\eta^{-1}D_{f}\big{(}\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big{)}$
2:  $\widehat{\pi}(a|s)$
In this subsection, we present an offline learning algorithm for $f$-divergence regularized contextual dueling bandit, $f$-CDB, in Algorithm 4. $f$-CDB first leverages maximum likelihood estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset. Then the algorithm constructs the output policy $\widehat{\pi}$ that maximizes the f-divergence regularized objective induced by ${\bar{g}}$. Similar to Algorithm 2, we do not require any pessimism in $f$-CDB. The following theorem provides an upper bound of Algorithm 4.
Theorem D.5.
For any sufficiently small $\epsilon\in(0,1)$, and $\eta,\alpha>0$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big{)}$ is sufficient to guarantee that the output policy $\widehat{\pi}$ of Algorithm 4 is $\epsilon$-optimal.
The following theorem provides a lower bound for offline $f$-divergence regularized contextual dueling bandit with strongly convex $f$.
Theorem D.6.
For any $\epsilon\in(0,1)$, $\alpha,\eta>0$, and offline RL algorithm Alg, there is an $\alpha$-strongly convex $f$ and $f$-divergence regularized contextual dueling bandit instance such that Alg requires at least $\Omega\big{(}\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big{)}$ samples to return an $\epsilon$-optimal policy.
Remark D.7.
Theorem D.6 indicates that, when $\epsilon$ is sufficiently small, to produce an $\epsilon$-optimal policy, any algorithm for offline $f$-regularized contextual bandits with strongly convex $f$ requires at least $\widetilde{\Omega}(\alpha^{-1}\eta\epsilon^{-1})$ samples. This lower bound matches the sample-complexity upper bound in Theorem D.5, indicating that Algorithm 4 is nearly optimal.
Appendix E Missing Proof from Appendix D
E.1 Proof of Theorem D.1
$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (E.1)
Here, $\Gamma_{n}$ is defined in (D.1). We abuse the notation and define $b(\cdot)$ as
$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (E.2)
where $\Phi_{b}(s,a)=\big{|}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{|}$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation ${\bar{g}}$ obtained in Line 1 of Algorithm 3, after adjusted by some bias function $b$, is close to the true function ${g^{*}}$. The following lemma shows that this event holds with high probability.
Lemma E.1.
For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.